Two Angles are supplementary. One angle is 20 degrees less than 4 times the other. Find the measures of the angles... Thank u so much :DDD
Answer by Edwin McCravy(20059)

Two Angles are supplementary. One angle is 20 degrees less than 4 times the other. Find the measures of the angles...
Let the other angle be x.

Other angle = x

>>...One angle is 20 degrees less than 4 times the other...<<
Replace the words "the other" by "x"

>>...One angle is 20 degrees less than 4 times x...<<
Replace the words "4 times x" by "4x"

>>...One angle is 20 degrees less than 4x...<<
The words "20 degrees less than" before "4x" tells us that we
are to start with 4x and make it 20 degrees less.  To make it
less means to subtract from it.  So we are to make 4x less by
subtracting 20 from it. 

So replace the words "20 degrees less than 4x" by "4x-20".

>>...One angle is 4x-20...<<
Replace the word "is" by "="

>>...One angle = 4x-20...<<
Now we have

>>...One angle = 4x-20...<<
and

Other angle = x

>>...Two Angles are supplementary...<<<
That means




(4x-20) + x = 180
4x - 20 + x = 180
    5x - 20 = 180
         5x = 200
          x = 40

So One angle = 4x-20 = 4(40)-20 = 160-20 = 140°

and

"the other angle" = x = 40° 

Edwin